OpenStudy (kimberly_pr):

The function f(x) = 467(5)x represents the growth of a ladybug population every year in a wooded area. Adrianne wants to manipulate the formula to an equivalent form that calculates every 3 months, not every year. Which function is correct for Adrianne's purposes?

OpenStudy (kimberly_pr):

The answer choices are f(x)=67(5)^x f(x)=467(5^12)x/12 f(x)=467(5and1/4th)^4x f(x)=4672(5)^x

OpenStudy (kimberly_pr):

@welshfella

OpenStudy (kimberly_pr):

@Mahoganie.Carson

OpenStudy (kimberly_pr):

@lovefootball @FortyTheRapper